oliviac1111:

Which equations are true for x = –2 and x = 2? Select two options x2 – 4 = 0 x2 = –4 3x2 + 12 = 0 4x2 = 16 2(x – 2)2 = 0

snowflake0531:

\[x=2 ~and~x=-2\]\[x^2-4=0 \\ x^2=-4\\3x^2+12=0\\4x^2=16\\2(x-2)^2=0\] Plug in 2 and -2 as x into each choice, and whichever 2 equals the right hand side of the equal sign, is the two you should choose .-.

oliviac1111:

thanks luv

snowflake0531:

yw~